OpenStudy (anonymous):

A cyclist rides her bike at a speed of 21 kilometers per hour. What is this speed in kilometers per minute? How many kilometers will the cyclist travel in 2 minutes?

OpenStudy (anonymous):

Speed = km/min Distance traveled in 2 minutes = km

OpenStudy (math&ing001):

\[21 \frac{ km }{ hour }=21\frac{ km }{ 60\min }=\frac{ 21 }{ 60 }km/\min\]